OpenStudy (anonymous):

y + 2 = -2(x + 3). how do you graph this ?

OpenStudy (anonymous):

this is the equation of a straight line just determine any two points of this line and join them ex: if x=0=>y+2=-2(0+3)=>y+2=-6=>y=-8 so plot (0;-8) chosse another point and join them

OpenStudy (anonymous):

put any two values of x... you will get 2 values of y.. join the two (x,y) coordinates formed

OpenStudy (anonymous):

can you show me ?

OpenStudy (anonymous):

put x = 0 .. y = -8 put x = 1 .. y = -10 we have two coordinates (0,-8) and (0,-10) Plot both points and join them by straight line..

OpenStudy (anonymous):



OpenStudy (anonymous):

@monica0597 @haganmc @Katiebae @rymdenbarn @Sunshine447 @amber517 @Sukhvir

OpenStudy (anonymous):

did i get it ?

OpenStudy (anonymous):

no:( the point is (0;-8) so on the y axis,8 units down

OpenStudy (anonymous):

so -4 -8 ?

OpenStudy (anonymous):

which two points did you use?

OpenStudy (anonymous):

-4 -8

OpenStudy (anonymous):

the coordinates must be written as (x;y) what you've written doesn't make much sense the 2 points for example can be (0;-8) and (-3;-2) you choose the x and find the corresponding y